OpenStudy (anonymous):

vertex form:\[y=x^2+10x+32\]

OpenStudy (anonymous):

after completing the square i keep getting \(y=(x-5)^2+7\) but the answer is \(y=(x+5)^2+7\)

hero (hero):

You completed the square wrong. No way should you get (x-5)

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

will*

hero (hero):

\[ y=x^2+10x+32 \\y - 32 = x^2 + 10 \\ y - 32 + 25 = x^2 + 10 + 25\\ y - 7 = (x+5)^2 \\ y = (x + 5)^2 + 7\]

OpenStudy (anonymous):

i think i get it now...thanks

hero (hero):

I know they teach these in a weird way. They try to show you how to complete the square without subtracting the 32 from both sides, but doing it that way is silly.

OpenStudy (anonymous):

yeah i know thanks though...i see it, i was skipping steps